OpenStudy (anonymous):

100 radians to degrees?

OpenStudy (anonymous):

which is...

OpenStudy (anonymous):

pi rad = 180 deg 1 rad = 180/pi deg 100 rad = 180*100/pi

OpenStudy (anonymous):

Follow what @rizwan_uet wrote down!

OpenStudy (anonymous):

but... 5729 doesn't make sense!

OpenStudy (anonymous):

it does make sense

OpenStudy (anonymous):

5729 degrees?

OpenStudy (anonymous):

terminal side is rotated this much degrees mean alot of revolutions

OpenStudy (anonymous):

wow.... okay... i guess thnx